# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems for Social Media

### v2

@@ 8.3.1. User Informedness
While social media serves many different purposes, CCS are specifically concerned with social media’s ability to inform the public. Such systems should thus be evaluated by how well they facilitate user informedness: for example, the degree to which they inform users about the accuracy of information, minimize the spread of misleading posts, or otherwise improve information quality on a platform. Our literature review surfaced ample evidence that the notes produced by CCS do have the power to increase user informedness, both by decreasing an individual’s belief in misleading information, and by decreasing the spread of this information (wojcikBirdwatchCrowdWisdom2022b; renaultCollaborativelyAddingContext2024a; slaughter25; chuaiCommunitybasedFactcheckingReduces2024a; rollout21; diffusion21). Additionally, there is evidence (kimDifferentialImpactIndividual2025e) that these notes may produce less of a backfiring effect than forms of social correction that is not mediated by a CCS (mosleh21). But while published notes have been shown to be effective at promoting understanding, we must also question how effective CCS are at producing these notes. Several of these same studies found that Twitter’s implementation of CCS produces notes too slowly to prevent the spread of the most viral content (rollout21; chuaiCommunitybasedFactcheckingReduces2024a; renaultCollaborativelyAddingContext2024a). There is also a significant gap in the research on the question of coverage, or how many of the misleading posts on a platform receive “helpful” notes, which is currently unanswered. The research suggests that there is a CCS design trade-off between speed/coverage and quality/helpfulness: X’s implementation has optimized for the latter, perhaps at the expense of the former. Whether this is the right decision should be up for debate.
@@ 8.3.2. Distribution of Power
One of the selling points of CCS is that they shift power over content moderation from the platform to the community. Giving users more of a say over the content that they see on their platform can make moderation decisions more legitimate and trusted (shusas24; liu23; drolsbachCommunityNotesIncrease2024a), which is an admirable goal. Yet such claims require further scrutiny, and CCS should be evaluated to the extent that they actually empower users and do not merely allow platforms to shirk responsibility for misleading content. Evidence from our literature review suggests that bridging-based algorithms, by design, fail to produce notes for the most polarizing content (bouchaudAlgorithmicResolutionCrowdsourced2025b). For example, X’s implementation of the bridging algorithm requires polarized contributors to agree that a note is helpful in order to publish it. The editorial function of the bridging algorithm can thus limit the type and scope of contributions in two important ways: in terms of framing, the algorithm may prefer centrist points of view and content that could be agreeable to both sides of an issue; in terms of selection, agreement on a note may not occur for contentious topics, preventing its publication. Whether or not users at the ends of the ideological spectrum should be empowered to make moderation decisions is a valid question, but it is uncontroversial that the most polarizing content needs moderation, too. Research has found evidence of harmful content being handled poorly by the system, either because potential notes never reach helpful status (chujyo25), or because notes that do are not actually helpful for certain types of content, such as disinformation that is couched in humor (matamoros25). Zooming out, the low publishing rate for suggested notes (around  10% on X (Wirtschafter_Majumder_2023) and  6% on Meta [^8] https://x.com/guyro/status/1965800300557857053) gives us pause that the system is actually empowering these note authors. A related, underexplored area for research involves studying which note contributors are empowered by CCS. This research area is difficult to study holistically, as these systems are pseudonymous, but some work has begun to explore it qualitatively (yoon25) and via partnerships with platforms (allen22). More research is needed to understand how CCS can empower users of certain demographics or political views over others.
@@
Our design space offers several ways that designers can further empower users via CCS. First and foremost is innovation in the participation aspect that would increase the set and diversity of contributors. The challenge for the platforms will be ensuing that quality remains as the participating contributor base increases, which is a tension that has been thoroughly studied in the crowdsourcing literature (danielQualityControlCrowdsourcing2018), for example. Innovation in the inputs aspect may increase the quality of notes across the board, perhaps by giving note authors helpful feedback (mohammadiAIFeedbackEnhances2025a), empowering more users to make successful contributions to the system. Another route to shift power to users is to disempower the platform. Though CCS are often described as ways to reduce platforms’ centrality in content moderation, platforms still exert power over the system in important ways, such as by deciding who is admitted as a contributor. Platforms could further guarantee their neutrality through changes to the transparency aspect, such as by making their contributor selection process public. But considering that X’s CCS implementation is the only even partially open system that we examined, a first step is for other platforms to follow X’s lead and open source their note curation algorithm and the crowdsourced data that they collect.
@@ 8.3.3. Fairness
CCS are often touted for their potential to implement content moderation in an unbiased way. Yet these systems are not inherently free from bias. Results from our literature review detail bias in contributor behavior, in terms of the types of accounts that they leave notes on, as well as how they rate suggested notes (allen22). Bridging algorithms have helped mitigate bias in the note curation process, with the caveats noted above, but additional bias can still be introduced—for example, in the note creation process, or in the selection of contributors. To give a concrete example: X’s bridging algorithm finds agreement among contributors, not the entire user base. If contributors’ beliefs are not representative of users’ beliefs then the algorithm’s outcomes may be a biased representation of the community, and not necessarily have broad appeal among users. There is some evidence that so far X’s implementation has not suffered from this problem, but it is worth considering. Also, accusations of bias in top-down moderation are often due to the opacity of such approaches and may be mitigated by transparency efforts. Only X’s implementation has provided any transparency, and yet the lack of transparency about the contributor base has made analyzing them for bias impossible. Finally, the approach of displaying context rather than removing posts is arguably fairer in that it lessens the damage from false positives, by empowering users to come to their own conclusions about content’s veracity (pengRescuingCounterspeechBridgingBased2024a). However, there is a trade off: leaving up questionable content, even with helpful context, can potentially do harm to those targeted by the content.
@@
Designers should consider several aspects from our design space that can influence the fairness properties of a CCS. First off, all platforms have taken the approach that it is fairer not to remove misleading content. However designers may explore other options from the platform treatment dimension to identify certain cases where the platform should take action on content flagged by the crowd as misleading. Second, current implementations tend to display a single, definitive note, even if there may be several that are determined to be “helpful”. The presentation aspect offers alternatives that could further the aim of allowing users to come to their own conclusions, such as by showing multiple notes when there are multiple with helpful ratings. Additionally showing information about the note author could be helpful, as it could help users identify the credibility of the source. This may be at odds with contributor privacy, which current implementations have opted to protect. Finally, innovations in the curation aspect could explore other methods of selecting notes to display that may be regarded as fairer. Current implementations of the bridging algorithm use past ratings to identify contributors with diverse perspectives: perhaps it would be fairer to instead use in-app behavior? This could provide a fuller picture of a contributor’s ideological leanings. Encouraging more transparency would also give assurances of fairness, both in how contributors are selected, and how notes are published.
@@ 8.4. Limitations
Since there was no agreed upon name for CCS before our paper, we decided to focus our literature review on studies of Twitter-style CCS, and thus may have missed earlier work that would have meaningfully informed the design space. Our analysis of platforms’ implementations of CCS is based on public documentation and may not perfectly reflect the actual features of the deployed systems. Finally, our design space includes several non-exhaustive, speculative options that were not seen in any implementations, but nonetheless are included to illuminate alternative possibilities.
@@ 9. Conclusion
Crowdsourced Context Systems are increasingly prevalent in the content moderation strategies of prominent social media platforms. To understand these systems, their impacts, and how they may be better designed, we conducted a systematic literature review of existing research studying implementations of CCS and analyzed several major social media platforms’ implementations of such systems. Despite recent trends towards the adoption of CCS, we emphasize that they are only one piece of a larger content moderation suite, and suggest against using them as drop-in replacements for traditional fact-checking approaches. The framework we produced can be used to reason about these systems, and consists of a theoretical model and a design space. We additionally discuss some of the key normative concerns that should be considered when designing CCS. Such systems have potential to address the challenge of misinformation, but to realize this potential, future design research from the HCI community must focus on ways to understand and improve these systems’ performance.

### v3

@@ 8.3.1. User Informedness
While social media serves many different purposes, CCS are specifically concerned with social media’s ability to inform the public. Such systems should thus be evaluated by how well they facilitate user informedness: for example, the degree to which they inform users about the accuracy of information, minimize the spread of misleading posts, or otherwise improve information quality on a platform. Our literature review surfaced ample evidence that the notes produced by CCS do have the power to increase user informedness, both by decreasing an individual’s belief in misleading information, and by decreasing the spread of this information (Wojcik et al., 2022; Renault et al., 2024; Slaughter et al., 2025; Chuai et al., 2024a, b; Drolsbach and Pröllochs, 2023). Additionally, there is evidence (Kim et al., 2025) that these notes may produce less of a backfiring effect than social correction that is not mediated by a CCS (Mosleh et al., 2021). But while published notes have been shown to be effective at promoting understanding, we must also question how effective CCS are at producing these notes. Several of these same studies found that Twitter’s implementation of CCS produces notes too slowly to prevent the spread of viral content (Chuai et al., 2024b, a; Renault et al., 2024). There is also a significant gap in the research on the question of coverage—how many of the misleading posts on a platform receive published notes—which currently remains unanswered. The research suggests that there is a CCS design trade-off between speed/coverage and quality/helpfulness: X’s implementation has optimized for the latter, perhaps at the expense of the former. Whether this is the right decision should be up for debate.
@@ 8.3.2. Distribution of Power
One of the selling points of CCS is that they shift power over content moderation from the platform to the community. Giving users more of a say over the content that they see on their platform can make moderation decisions more legitimate and trusted (Shusas and Forte, 2024; Liu et al., 2023; Drolsbach et al., 2024). Yet such claims require further scrutiny, and CCS should be evaluated to the extent that they actually empower users and do not merely allow platforms to shirk responsibility for misleading content. Evidence from our literature review suggests that bridging algorithms, by design, fail to produce notes for the most polarizing content (Bouchaud and Ramaciotti, 2025). For example, X’s implementation of the bridging algorithm requires polarized contributors to agree that a note is helpful in order to publish it. The editorial function of the bridging algorithm can thus limit the type and scope of contributions in two important ways: in terms of framing, the algorithm may prefer centrist points of view and content that could be agreeable to both sides of an issue; in terms of selection, agreement on a note may not occur for contentious topics, preventing its publication. Whether or not users at the ends of the ideological spectrum should be empowered to make moderation decisions is a valid question, but it is uncontroversial that the most polarizing content needs moderation, too. Research has found evidence of harmful content being handled poorly by the system, either because potential notes never reach helpful status (Chujyo et al., 2025), or because notes that do are not actually helpful for certain types of content, such as disinformation that is couched in humor (Matamoros-Fernández and Jude, 2025). Zooming out, the low publishing rate for suggested notes (around  10% on X (Wirtschafter and Majumder, 2023) and  6% on Meta [^9] https://x.com/guyro/status/1965800300557857053) gives us pause that these systems are actually empowering note authors. A related, underexplored area for research involves studying which note contributors are empowered by CCS. This research area is difficult to study holistically, as these systems are pseudonymous, but some work has begun to explore it qualitatively (Yoon et al., 2025) and via partnerships with platforms (Allen et al., 2022). More research is needed to understand how CCS may empower users of certain demographics or political views over others.
@@
Our design space offers several ways that designers can further empower users via CCS. First and foremost is innovation in the Participation aspect that would increase the set and diversity of contributors. The challenge for the platforms will be ensuing that quality remains as the contributor base increases, which is a tension that has been thoroughly studied in the crowdsourcing literature (Daniel et al., 2018), for example. Innovation in the Inputs aspect may increase the quality of notes across the board, perhaps by giving note authors helpful feedback (Mohammadi and Yasseri, 2025) that empowers more users to make successful contributions. Another route to shift power to users is to disempower the platform. Though CCS are often described as ways to reduce platforms’ centrality in content moderation, platforms still exert power over the system in important ways, such as by deciding who is admitted as a contributor. Platforms could further guarantee their neutrality through changes to the Transparency aspect, such as by making their contributor selection process public. But considering that X’s CCS implementation is the only even partially open system that we examined, a first step is for other platforms to follow X’s lead and open source their note-curation algorithm and the crowdsourced data that they collect.
@@ 8.3.3. Fairness
CCS are often touted for their potential to implement content moderation in an unbiased way. Yet these systems are not inherently free from bias. Results from our literature review show bias in contributor behavior, in terms of the types of accounts that they leave notes on, as well as how they rate suggested notes (Allen et al., 2022). Bridging algorithms have helped mitigate bias in the note selection process, with the caveats noted above, but additional bias can still be introduced—for example, in the note creation process, or in the selection of contributors. To give a concrete example: X’s bridging algorithm finds agreement among contributors, not the entire user base. If contributors’ beliefs are not representative of users’ beliefs then the algorithm’s outcomes may be a biased representation of the community, and not necessarily have broad appeal among users. There is some evidence that so far X’s implementation has not suffered from this problem, but it is worth considering. Also, accusations of bias in top-down moderation are often due to the opacity of such approaches and may be mitigated by transparency efforts. Only X’s implementation has provided any transparency, and yet the lack of transparency about the contributor base has made analyzing it for bias impossible. Finally, the approach of displaying context rather than removing posts is arguably fairer in that it lessens the damage from false positives, by empowering users to come to their own conclusions about content’s veracity (Peng and Grimmelmann, 2024). However, there is a trade-off: leaving up questionable content, even with helpful context, can potentially do harm to those targeted by the content.
@@
Designers should consider several aspects from our design space that can influence the fairness properties of a CCS. First off, all platforms have taken the approach that it is fairer not to remove misleading content. However, designers may explore other options from the Platform Treatment dimension to identify certain cases where the platform should take action on content flagged by the crowd as misleading. Second, current implementations tend to display a single, definitive note, even if there may be several that are determined to be “helpful”. The Presentation aspect offers alternatives that could further the aim of allowing users to come to their own conclusions, such as by showing multiple notes when there are multiple with helpful ratings. Additionally showing information about the note author could be helpful, as it could help users identify the credibility of the source. At the same time, this feature may be at odds with contributor privacy, which current implementations have opted to protect. Finally, innovations in the Curation aspect could explore other methods of selecting notes to display that may be regarded as fairer. Current bridging algorithm implementations use past ratings to identify contributors with diverse perspectives: perhaps it would be fairer to instead use in-app behavior? This could provide a fuller picture of a contributor’s ideological leanings. Encouraging more Transparency would also give assurances of fairness, both in how contributors are selected, and how notes are published.
@@ 8.4. Limitations
Our work here may have several limitations. Since there was no agreed upon name for CCS before our paper, we decided to focus our literature review on studies of Twitter-style CCS, and thus may have missed earlier work that would have meaningfully informed the design space. Our analysis of platforms’ implementations of CCS is primarily based on public documentation from a specific point in time and may not perfectly reflect the actual features of the deployed systems.
@@ 9. Conclusion
Crowdsourced Context Systems are increasingly prevalent in the content moderation strategies of prominent social media platforms. To understand these systems, their impacts, and potential design improvements, we conducted a systematic literature review of CCS research and analyzed several major platforms’ implementations. Despite recent trends towards the adoption of CCS, we emphasize that they are only one piece of a larger content moderation suite, and advise against using them as drop-in replacements for traditional fact-checking approaches. The framework we produced can be used to reason about these systems, and consists of a theoretical model and a design space. We additionally discuss some of the key normative concerns that should be considered when designing CCS. Such systems have potential to address the challenge of misinformation, but to realize this potential, future design research from the HCI community must focus on ways to understand and improve these systems’ performance.
Acknowledgements.
This material is partially based upon work supported by the National Science Foundation under CISE Graduate Fellowship Grant No. 2313998 and SaTC Grant No. 2120651, as well as the John D. and Catherine T. MacArthur Foundation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the John D. and Catherine T. MacArthur Foundation or the National Science Foundation. Thank you to James Grimmelmann and Kenny Peng for feedback on earlier versions of this work.